# Robust Identity Perceptual Watermark Against Deepfake Face Swapping

### v2

@@ 4.1 Implementation Details
4.2 Experiment on CelebA-HQ
In this section, experiments are conducted on the testing set of CelebA-HQ. Since no existing work has performed detection and source tracing on face swapping images using a single watermark to our knowledge, we selected the robust watermarking frameworks (HiDDeN [61], MBRS [18], and CIN [28]), proactive robust Deepfake watermarking frameworks (RDA [58], SepMark [51], and ARWGAN [16]), proactive semi-fragile Deepfake watermarking frameworks (FaceSigns [31]), and proactive Deepfake distortion frameworks (AntiForgery [44] and CMUA [15]) for comparison. For fairness, the publicly available trained model weights that correspond to optimal performance are directly adopted. Statistics of the recent proactive separable watermark framework, SepMark [^1] SepMark [51] is only compared in Table 1 because the reported results are on CelebA-HQ, but the parameters are unrevealed for the common manipulations, leading to unavailability in maintaining a fair comparison in the remaining experiments. [51], are copied from the published paper due to unavailable source code.
4.2.1 Qualitative Evaluation
We visualized randomly drafted images to validate the visual quality of the watermarked images. As exhibited in Fig. 2, despite knowing the existence of the watermarks, differences between the raw and watermarked images in the first and second columns, respectively, are visually imperceptible. Starting from column 3, we illustrated the visual effects of the common manipulations on the watermarked images, one for each column. In Fig. 3, the face swapping effects are exhibited in every two columns starting from column 4. The source images that provide the desired synthetic facial identities for each row are placed in the first column. Columns 2 and 3 display the raw and watermarked target images, while every pair of the remaining columns illustrates the face swapping results on the raw and watermarked target images, respectively. It is worth noting that no distortion can be observed in the face swapping results using watermarked target images. In other words, the encoded watermarks do not affect the visual qualities of the synthetic faces, preserving the possible benign usages of Deepfake face swapping in the industry.
4.2.2 Quantitative Evaluation
Visual Quality. We employed the average peak signal-to-noise ratio (PSNR) and structural similarity index measure (SSIM) [49] between the raw and watermarked images to evaluate the visual quality of watermarked images. As Table 1 reports, while the goal of embedding watermarks into raw images imperceivably stands for all contrastive methods, our proposed approach achieves the state-of-the-art visual quality in terms of PSNR and SSIM at both resolution levels. Meanwhile, it can be observed that although pursuing the same objective in visual quality, the proactive distortion models, AntiForgery [44] and CMUA [15], have revealed relatively unsatisfactory performance since the noises and perturbations are directly added to raw images while watermarking approaches execute image reconstructions.
@@ 4.2.2 Quantitative Evaluation
Table 4: Quantitative comparison on LFW at 128 and 256 resolutions regarding the bit-wise recovery accuracy of the watermarks under common manipulations and malicious face swapping manipulations. The best performance for each resolution is in bold and the second best performance is marked with underscore.
@@
Manipulations HiDDeN  [61] $128\times 128$ MBRS  [18] $128\times 128$ RDA  [58] $128\times 128$ CIN  [28] $128\times 128$ ARWGAN  [16] $128\times 128$ Ours $128\times 128$ MBRS  [18] $128\times 128$ FaceSigns  [31] $128\times 128$ Ours $256\times 256$
Dropout 82.84% 99.99% 89.72% 100.00% 97.28% 99.96% 99.30% 56.45% 99.97%
Resize 81.97% 100.00% 99.89% 99.98% 95.07% 99.99% 100.00% 98.91% 98.16%
Jpeg 56.17% 99.42% 66.77% 97.06% 70.86% 99.38% 99.08% 75.64% 99.99%
GaussianNoise 51.30% 99.80% 60.83% 87.54% 53.11% 99.63% 55.61% 59.90% 99.92%
SaltPepper 52.93% 99.63% 67.31% 91.71% 62.96% 95.89% 61.15% 67.99% 95.99%
GaussianBlur 72.89% 100.00% 99.88% 99.99% 86.39% 99.99% 73.93% 99.01% 97.11%
MedBlur 82.35% 99.99% 99.98% 99.98% 96.93% 99.99% 99.38% 99.91% 98.13%
Average 68.63% 99.83% 83.48% 96.61% 80.37% 99.26% 84.06% 79.69% 98.47%
SimSwap [5] 49.96% 49.92% 50.04% 50.25% 50.00% 99.37% 50.04% 49.77% 99.96%
InfoSwap [10] 50.06% 50.04% 50.10% 50.95% 50.86% 98.69% 50.23% 49.91% 98.67%
UniFace [52] 53.74% 49.78% 70.37% 49.92% 59.16% 98.48% 50.18% 52.25% 97.06%
E4S [27] 49.34% 50.08% 67.78% 50.22% 48.64% 97.27% 49.99% 50.15% 97.00%
Average 50.77% 49.96% 59.57% 50.33% 52.18% 98.45% 50.11% 50.52% 98.17%
where $d_{\textrm{ham}}$ computes the hamming distance [43] between the recovered and original watermarks $m_{\textrm{rec}}$ and $m$, and $l$ refers to the watermark length. In the experiment, each type of manipulation in $P_{\textrm{common}}$ and $P_{\textrm{swap}}$ is performed on the watermarked image $I_{\textrm{rec}}$ before extracting the watermark via the decoder.
The watermark recovery performance in the comparative experiment is reported in Table 2. Specifically, considering the common manipulations, our method outperforms HiDDeN [61], RDA [58], CIN [28], and ARWGAN [16] at the $128\times 128$ resolution with a better average accuracy, and our performance is competitive against MBRS [18] with an acceptable average accuracy above $99\%$, achieving the second best statistic. However, it can be observed from Table 1 that MBRS heavily suffers visual quality issues with unsatisfactory PSNR and SSIM values ($33.01$ and $0.775$) in order to achieve the best bit-wise watermark recovery accuracy in Table 2. As for the $256\times 256$ resolution, we reached an average accuracy of $99.07\%$ on the common manipulations, outperforming the existing state-of-the-art methods. On the contrary, although accomplishes reasonable visual quality as listed in Table 1, MBRS experiences a large performance damping on the common manipulations in Table 2 with respect to the watermark recovery accuracy compared to the $128\times 128$ resolution version.
Watermark Recovery on Face Swapping. In this paper, the face swapping manipulation pool contains merely SimSwap [5] in the training phase, and we validated the watermark recovery performance for both in- and cross-manipulation settings by introducing other recent state-of-the-art face swapping algorithms, InfoSwap [10], UniFace [52], and E4S [27], in the testing phase. As shown in Table 2, our model consistently recovers the desired watermark messages at average accuracies of 98.02% and 96.98% for $128\times 128$ and $256\times 256$ resolutions, respectively, while HiDDeN, MBRS, CIN, and ARWGAN all obtain accuracies around 50% and RDA has accuracies below 75% against all synthetic models. Moreover, it can be concluded from the cross-manipulation tests on unseen synthetic models that by simply tuning with SimSwap, the hidden watermarks are pleasantly maintained by our method since the goals for different face swapping algorithms are identical. Meanwhile, the semi-fragile watermarking framework, FaceSigns [31], although obtains low watermark recovery accuracies as designed when facing synthetic manipulations that are expected to destroy the embedded watermarks, its performance regarding the common manipulations is unexpectedly unsatisfactory besides illustrating low visual quality as reported in Table 1, leading to an average recovery accuracy of $80.39\%$.
Discussion. In general, Salt and Pepper and Gaussian Noise are the most challenging common manipulations that cause fluctuations in the watermark recovery accuracies for all watermarking algorithms as summarized in Table 2. This is consistent with the visualization results in Fig. 2 (col. 6 & 9) such that these two manipulations add a plethora of perceivable noises to the images and thus heavily jeopardize the hidden watermarks. On the other hand, the contrastive robust watermarking frameworks fail to maintain robustness when facing the state-of-the-art face swapping models. As a result, our proposed model is the only one that maintains the bit-wise watermark recovery accuracies above 95% for all common manipulations and above 90% for all face swapping manipulations, while the contrastive methods are heavily affected by at least one of the manipulations.
@@
Table 5: Deepfake detection performance in AUC scores against different face swapping algorithms on CelebA-HQ at $128\times 128$ and $256\times 256$ resolutions. The last row evaluates the detection performance on a mixture of all face swapping algorithms. The best performance at each resolution is in bold and the second best performance is marked with underscore.
@@
SimSwap [5] 39.73% 71.15% 75.30% 88.94% 60.37% 69.01% 55.91% 87.66% 98.49% 98.19%
InfoSwap [10] 60.82% 65.50% 85.11% 80.50% 55.51% 52.13% 48.29% 61.39% 99.02% 99.52%
UniFace [52] 71.49% 70.34% 72.45% 79.41% 61.58% 67.35% 82.16% 82.73% 99.99% 98.99%
E4S [27] 43.40% 53.70% 63.63% 61.05% 60.88% 47.19% 64.93% 73.13% 97.71% 98.70%
Mixed 49.71% 63.04% 75.96% 80.05% 62.84% 57.13% 64.41% 72.52% 98.66% 98.97%
FF++ [38] 97.60% 90.50% 96.81% 95.57%
4.3 Cross-Dataset Experiment
We further evaluated our watermarking framework on an unseen dataset, Labeled Faces in the Wild (LFW) [14], regarding the generalization ability. Performance evaluation with respect to the visual quality is listed in Table 3. Similar results as in Table 1 can be summarized even on a different dataset. In particular, for the $128\times 128$ resolution, our approach consistently exhibits substantial performance advantages over the comparative ones. As for the $256\times 256$ resolution, MBRS [18] and our approach both have executed watermarking with promising visual qualities while FaceSigns [31] fails to do so.
Meanwhile, as listed in Table 4, the challenging manipulations concluded in Section 4.2 consistently lead to unsatisfactory bit-wise watermark recovery accuracies in this section. Consequently, although slightly affected by Salt and Pepper ($95.89\%$ and $95.99\%$) and Gaussian Blur ($99.99\%$ and $97.11\%$) at $128\times 128$ and $256\times 256$ resolutions, respectively, we successfully maintained state-of-the-art accuracies above $95\%$ for all common and face swapping manipulations.
Furthermore, obvious accuracy declines of our approach on UniFace [52] at the $256\times 256$ resolution and on E4S [27] at both resolutions can be observed in Table 2 and Table 4 on CelebA-HQ and LFW. Besides the challenges caused by the cross-manipulation setting, the former is mainly due to the setting of accepting only half-precision floating-point inputs by UniFace. Although the synthetic pipeline is unaffected, this can lead to discrepancies in the watermark recovery process at higher resolution levels because of information loss. On the other hand, the performance on E4S is affected by adopting an extra color channel, alpha, during image synthesis, which reduces the watermark feature density in the face swapped images, leading to a relatively lower watermark recovery accuracy.
4.4 Watermark Collision Check
As introduced in Section 3.2, the binary identity perceptual watermarks are derived from identity embeddings generated via well-designed face recognition tools [^2] https://github.com/serengil/deepface. To validate the claim that the identity perceptual watermarks retain the collision-free characteristic of the original identity embeddings, we performed collision checks at different watermark lengths on CelebA-HQ [19] and LFW [14]. In specific, for the sets of watermarks $M_{i}$ and $M_{j}$ of a pair of arbitrary identities $i$ and $j$, the following statement stands,
$\forall x\in M_{i}\forall y\in M_{j}(x\neq y).$ (11)
@@ 4.4 Watermark Collision Check
Table 6: Robustness ablation test on common manipulations for easier and harder parameter values that are unseen during training. The bit-wise watermark recovery accuracies are reported. Long common manipulation names are abbreviated.
@@
In this paper, images from CelebA-HQ with 6,217 unique identities are adopted to establish $\phi$ and produce watermarks following the pipeline in Section 3.2. Thereafter, the watermarks of images in LFW with 5,749 unseen identities are derived following the fixed rule accordingly.
As a result, there is no collision in the experiment for CelebA-HQ and LFW at both $128\times 128$ and $256\times 256$ resolutions in our collision check. Meanwhile, upon passing the collision check even though not seeing the identities in LFW when defining $\phi$, it can be concluded that the number of unique identities in CelebA-HQ is sufficient to avoid the potential bias, and the cross-dataset performance is therefore preserved.
4.5 Deepfake Detection
In this section, the Deepfake detection task is proactively achieved by evaluating the recovered watermark $m_{\textrm{rec}}$. In specific, for a target image $I$ with identity $t$ to be proactively protected with identity perceptual watermark $m$, the watermarked image $I_{\textrm{rec}}$ is firstly derived via the encoder of our framework. Then, when a face swapping algorithm from $P_{\textrm{swap}}$ is applied on $I_{\textrm{rec}}$ to generate $I_{\textrm{fake}}$ with identity $s$, the identity perceptual watermark $m_{s}$ is derived based on the image content of $I_{\textrm{fake}}$, and the robust watermark is also recovered, denoted as $m_{\textrm{rec}}$, via the decoder of our framework. In the end, the two watermarks, $m_{s}$ and $m_{\textrm{rec}}$, are compared bit-wisely and the falsification can be addressed based on their matching rate.
Since the comparative proactive watermarking approaches do not assign semantics to their watermarks, Deepfake detection cannot be accomplished without knowing the matched watermark for each image in advance, and this is usually unavailable in real-life scenarios. Meanwhile, the unsatisfactory watermark recovery accuracies in Section 4.2 suggest that the detection performance evaluation can be trivial for them. Therefore, in this experiment, we adopted several state-of-the-art passive Deepfake detectors in recent years for comparison, namely, SBIs [39], RECCE [4], and CADDM [7]. We also consider the popular milestone detector, Xception [^3] The well-trained model weights are adopted from a recent benchmark study [53].  [38], as one of the contrastive algorithms. All the models are utilized with their published weights corresponding to the optimal performance.
Since the accuracy metric requires a fixed threshold value which can lead to biases due to data distributions, we employed the area under the receiver operating characteristic (ROC) curve (AUC) score to evaluate the detection performance based on the matching rate. In a nutshell, a high AUC score demonstrates the probability that a random positive sample scores higher than a random negative sample from the testing set, in other words, the ability of the classifier to distinguish between real and fake samples. The experiment is conducted on CelebA-HQ at $128\times 128$ and $256\times 256$ resolutions by testing against each face swapping model in $P_{\textrm{swap}}$ and a mixture [^4] When conducting this experiment, samples are equally and randomly drawn for each synthetic algorithm to match up the same total quantity as the real ones. of them. To establish a balanced ratio for real and fake samples, the real images in CelebA-HQ are leveraged after randomly applying common manipulations from $P_{\textrm{common}}$.
The detection performance in AUC scores is demonstrated in Table 5. It can be observed that, although exhibit reasonable detection performance in lab-controlled scenarios, in this experiment, the passive detectors generally come up with poor AUC scores of at most 88.94%. Moreover, various statistics are observed below 70% regarding each passive model, demonstrating substantial fluctuations in the detection performance. We also listed their performance on the training dataset, FaceForensics++ [38] (FF++) dataset, for a reference. As a result, huge performance damping can be concluded for the passive detectors since the face swapping models are unseen and can generate faces with imperceptible detectable artifacts. Meanwhile, the passive detectors trained in lab-controlled scenarios fail to stay robust when facing common manipulations on the real samples that simulate the real-life scenarios.
On the contrary, our proposed approach successfully retains promising AUC scores above 97% for both seen and unseen face swapping manipulations at both resolutions. This implies that the matching rates between $m_{s}$ and $m_{\textrm{rec}}$ for images after common manipulations (real) are consistently higher than those for images after face swapping manipulations (fake). In other words, our identity perceptual watermarking framework not only robustly survives via image manipulations, but favorably distinguishes real and fake samples with respect to the changes in facial identities. As a result, for a mixture of all four manipulations including three unseen ones, our approach retains satisfactory detection performance at 98.66% and 98.97% at $128\times 128$ and $256\times 256$ resolutions, respectively, demonstrating superior generalization ability.
4.6 Ablation Study on Common Manipulations

### v3

@@ IV-B2 Quantitative Evaluation
TABLE V: Quantitative comparison on LFW at the $128\times 128$ resolution regarding the bit-wise recovery accuracy of the watermarks under common manipulations and malicious face swapping manipulations. The best performance for each resolution is in bold and the second best performance is marked with underscore.
@@
Manipulations HiDDeN [34] MBRS [35] RDA [40] CIN [36] IMD [46] MaLP [47] ARWGAN [39] SepMark [21] WEvade [69] Ours
Dropout 82.84% 99.99% 89.72% 99.99% 0.5833 0.4785 97.28% 99.98% 76.81% 99.96%
Resize 81.97% 99.99% 99.89% 99.98% 0.4039 0.4828 95.07% 99.93% 68.59% 99.99%
Jpeg 56.17% 99.42% 66.77% 97.06% 0.5088 0.1273 70.86% 83.42% 68.63% 99.38%
GaussianNoise 51.30% 99.80% 60.83% 87.54% 0.4453 0.4704 53.11% 81.90% 68.98% 99.63%
SaltPepper 52.93% 99.63% 67.31% 91.71% 0.3789 0.2944 62.96% 99.78% 56.93% 95.89%
GaussianBlur 72.89% 99.99% 99.88% 99.99% 0.3404 0.4877 86.39% 91.97% 63.04% 99.99%
MedBlur 82.35% 99.99% 99.98% 99.98% 0.4082 0.4802 96.93% 89.72% 69.99% 99.99%
Average 68.63% 99.83% 83.48% 96.61% 0.4384 0.4030 80.37% 92.39% 67.57% 99.26%
SimSwap [60] 49.96% 49.92% 50.04% 50.25% 0.0024 0.7341 50.00% 57.44% 59.12% 99.37%
InfoSwap [61] 50.06% 50.04% 50.10% 50.95% 0.5258 0.1417 50.86% 73.98% 77.93% 98.69%
UniFace [62] 53.74% 49.78% 70.37% 49.92% 0.4024 0.4944 59.16% 63.10% 51.52% 98.48%
E4S [63] 49.34% 50.08% 67.78% 50.22% 0.0064 0.6432 48.64% 88.14% 71.03% 97.27%
DiffSwap [64] 42.36% 43.35% 61.94% 47.15% 0.5763 0.5616 59.25% 74.42% 75.12% 83.29%
FuseAnyPart [65] 72.26% 89.84% 28.22% 76.67% 0.4635 0.6390 67.16% 90.46% 86.94% 93.15%
Face-Adapter [66] 75.65% 85.01% 49.67% 75.18% 0.4542 0.6401 76.02% 77.74% 75.58% 88.91%
REFace [67] 48.85% 87.05% 49.10% 72.94% 0.4685 0.5668 70.60% 83.83% 64.96% 84.96%
Average 55.28% 63.13% 53.40% 59.16% 0.3624 0.5526 60.21% 76.14% 70.28% 93.02%
  • $\dagger$ IMD and MaLP are specially evaluated via cosine similarity according to the original paper. Higher values denote better watermark robustness.
TABLE VI: Quantitative comparison on LFW at the $256\times 256$ resolution regarding the bit-wise recovery accuracy of the watermarks under common manipulations and malicious face swapping manipulations. The best performance for each resolution is in bold and the second best performance is marked with underscore.
Manipulations MBRS [35] FaceSigns [15] SepMark [21] InvisMark [70] Ours
Dropout 99.30% 56.45% 99.99% 99.86% 99.97%
Resize 99.99% 98.91% 99.99% 97.09% 98.16%
Jpeg 99.08% 75.64% 83.41% 52.93% 99.99%
GaussianNoise 55.61% 59.90% 81.89% 66.51% 99.92%
SaltPepper 61.15% 67.99% 99.99% 56.29% 95.99%
GaussianBlur 73.93% 99.01% 91.96% 99.99% 97.11%
MedBlur 99.38% 99.91% 89.71% 99.99% 98.13%
Average 84.06% 79.69% 92.42% 81.81% 98.47%
SimSwap [60] 50.04% 49.77% 52.94% 52.86% 99.96%
InfoSwap [61] 50.23% 49.91% 80.43% 62.63% 98.67%
UniFace [62] 50.18% 52.25% 54.10% 53.01% 97.06%
E4S [63] 49.99% 50.15% 86.58% 74.15% 97.00%
DiffSwap [64] 77.11% 53.26% 77.38% 52.88% 98.59%
FuseAnyPart [65] 83.01% 49.13% 95.46% 53.63% 99.20%
Face-Adapter [66] 83.94% 50.63% 83.28% 53.11% 93.81%
REFace [67] 69.11% 50.17% 70.45% 52.87% 85.15%
Average 64.20% 50.66% 75.08% 56.89% 96.18%
IV-C Cross-Dataset Experiment
We further evaluated our watermarking framework on an unseen dataset, Labeled Faces in the Wild (LFW) [59], regarding the generalization ability. Performance evaluation with respect to the visual quality is listed in Table IV. Similar results as in Table I can be summarized even on a different dataset. In particular, for the $128\times 128$ resolution, our approach consistently exhibits substantial performance advantages over the comparative ones. As for the $256\times 256$ resolution, MBRS [35], InvisMark [70], and our approach have executed watermarking with promising visual qualities, while FaceSigns [15] and SepMark [21] fail to do so.
Meanwhile, as listed in Table V and Table VI, the challenging common manipulations concluded in Section IV-B consistently lead to unsatisfactory bit-wise watermark recovery accuracies in this section. Subsequently, although slightly affected by Salt and Pepper ($95.89\%$ and $95.99\%$) and Gaussian Blur ($99.99\%$ and $97.11\%$) at $128\times 128$ and $256\times 256$ resolutions, respectively, we successfully maintained state-of-the-art watermark recovery accuracies above $95\%$ for all common manipulations. As for watermark robustness against face swapping models, our approach generally performs consistently compared to that on CelebA-HQ, although slight accuracy drops can be observed on average in this cross-dataset scenario. Similarly, a higher average watermark recovery accuracy is achieved at the $256\times 256$ resolution than at $128\times 128$, which is mainly reflected in the diffusion-based models.
IV-D Watermark Collision Check
As introduced in Section III-B, the binary identity perceptual watermarks are derived from identity embeddings generated via well-designed face recognition tools [^3] https://github.com/serengil/deepface. To validate the claim that the identity perceptual watermarks retain the collision-free characteristic of the original identity embeddings, we performed collision checks at different watermark lengths on CelebA-HQ [57] and LFW [59]. In specific, for the sets of watermarks $M_{i}$ and $M_{j}$ of a pair of arbitrary identities $i$ and $j$, the following statement stands,
$\forall x\in M_{i}\forall y\in M_{j}(x\neq y).$ (11)
In this paper, images from CelebA-HQ with 6,217 unique identities are adopted to establish $\phi$ and produce watermarks following the pipeline in Section III-B. Thereafter, the watermarks of images in LFW with 5,749 unseen identities are derived following the fixed rule accordingly. As a result, there is no collision in the experiment for CelebA-HQ and LFW at both $128\times 128$ and $256\times 256$ resolutions in our collision check. Meanwhile, upon passing the collision check even though not seeing the identities in LFW when defining $\phi$, it can be concluded that the number of unique identities in CelebA-HQ is sufficient to avoid the potential bias, and the cross-dataset performance is therefore reliably preserved.
IV-E Deepfake Detection
In this section, the Deepfake detection task is proactively achieved by evaluating the recovered watermark $m_{\textrm{rec}}$. Specifically, for a target image $I$ with identity $t$ to be proactively protected with an identity perceptual watermark $m$, the watermarked image $I_{\textrm{rec}}$ is firstly derived via the encoder of our framework. Then, when a face swapping algorithm is applied on $I_{\textrm{rec}}$ to generate $I_{\textrm{fake}}$ with a new identity $s$, an identity perceptual watermark $m_{s}$ is then derived based on the image content of $I_{\textrm{fake}}$. Meanwhile, the robust watermark is also recovered, denoted as $m_{\textrm{rec}}$, via the decoder of our framework. In the end, the two watermarks, $m_{s}$ and $m_{\textrm{rec}}$, are compared bit-wisely and the falsification can be addressed based on their matching rate. In other words, since $m_{\textrm{rec}}$ represents a perfect recovery of the original embedded watermark that corresponds to the original facial identity and $m_{s}$ represents the watermark that corresponds to the identity in the current image content, a low similarity between the two implies distinct facial identities for images $I_{\textrm{rec}}$ and $I_{\textrm{fake}}$, indicating that Deepfake face swapping happens from $I_{\textrm{rec}}$ to $I_{\textrm{fake}}$. Contrarily, a promisingly high similarity between $m_{s}$ and $m_{\textrm{rec}}$ implies that the facial identity remains unchanged. It is worth noting that the entire proactive Deepfake detection via our proposed identity perceptual watermarks does not require any ground-truth watermarks.
@@ IV-E Deepfake Detection
TABLE VII: Deepfake detection performance in AUC scores against different face swapping algorithms on CelebA-HQ at $128\times 128$ and $256\times 256$ resolutions. The last row evaluates the detection performance on a mixture of all face swapping algorithms. The best performance at each resolution is in bold.
@@
SimSwap [60] 98.78% 99.99% 85.38% 98.63% 99.55% 99.98% 99.98% 99.98% 98.49% 98.19%
InfoSwap [61] 33.93% 48.98% 61.59% 81.55% 43.06% 56.32% 45.92% 56.56% 99.02% 99.52%
UniFace [62] 55.29% 46.03% 84.71% 81.64% 78.57% 53.79% 63.55% 64.61% 99.99% 98.99%
E4S [63] 30.13% 39.84% 55.64% 67.06% 26.83% 40.81% 28.63% 36.85% 97.71% 98.70%
DiffSwap [64] 65.40% 53.62% 56.71% 68.65% 86.18% 63.60% 81.93% 73.33% 92.73% 98.09%
FuseAnyPart [65] 31.15% 62.50% 77.42% 79.58% 74.71% 63.80% 74.23% 81.34% 93.96% 98.26%
Face-Adapter [66] 30.57% 62.52% 75.72% 66.64% 73.61% 75.51% 72.86% 67.83% 95.80% 99.15%
REFace [67] 29.49% 31.25% 66.51% 76.88% 65.32% 64.12% 65.77% 65.21% 82.49% 85.93%
Mixed 33.01% 66.89% 67.85% 75.19% 66.61% 69.94% 75.90% 69.25% 95.39% 97.57%
Since existing comparative proactive watermarking approaches do not assign semantics to their watermarks, Deepfake detection cannot be accomplished without knowing the matched watermark for each image in advance, and this is usually unavailable in real-life scenarios. Meanwhile, the unsatisfactory watermark recovery accuracies in Section IV-B suggest that evaluating detection performance can be trivial for them. Therefore, in this experiment, we adopted several popular and state-of-the-art passive Deepfake detectors in recent years for comparison, namely, Xception [2], SBIs [74], RECCE [75], and CADDM [76]. Since our approach is trained on CelebA-HQ with SimSwap accessible, we trained the detectors on CelebA-HQ against SimSwap, rather than directly adopting the model weights trained on other datasets, to guarantee fairness in the experiment. In particular, images in CelebA-HQ are recognized as real samples, and the fake samples are derived by executing face swapping on CelebA-HQ via SimSwap.
Since the accuracy metric requires a fixed threshold value, which can lead to biases due to data distributions, we employed the area under the receiver operating characteristic (ROC) curve (AUC) score to evaluate the detection performance based on the matching rate. In a nutshell, a high AUC score demonstrates the probability that a random positive sample scores higher than a random negative sample from the testing set, in other words, the ability of the classifier to distinguish between real and fake samples. The experiment is conducted on CelebA-HQ at $128\times 128$ and $256\times 256$ resolutions by testing against each face swapping model in $P_{\textrm{swap}}$ and a mixture [^4] When conducting this experiment, samples are equally and randomly drawn for each synthetic algorithm to match the same total quantity as the real ones. of them. To establish a balanced ratio for real and fake samples, the real images in CelebA-HQ are leveraged after randomly applying common manipulations from $P_{\textrm{common}}$.
The detection performance in AUC scores is demonstrated in Table VII. It can be observed that, except SBIs that reconstructs its own fake training samples following the designed pipeline, all models have achieved outstanding AUC scores on detecting SimSwap, where CADDM reaches 99.98% at the $128\times 128$ resolution and Xception achieves 99.99% at the $256\times 256$ resolution, slightly outperforming our promising performance of 98.49% and 98.19% at the two resolutions, respectively. This demonstrates the reliable detection ability of the passive detectors on seen manipulations. However, when facing unseen face swapping manipulations, the passive detectors generally come up with poor statistics of at most 86.18%. At the same time, many statistics are observed below 50% regarding the passive detectors, demonstrating substantial fluctuations in the detection performance against different synthetic models. For instance, RECCE achieves an 86.18% AUC score against DiffSwap at the $128\times 128$ resolution, but a poor 26.83% statistic is derived against E4S. As a result, huge performance damping can be concluded for the passive detectors since the face swapping models are unseen and can generate faces with imperceptible detectable artifacts.
Contrarily, our proposed approach successfully retains promising AUC scores above $90\%$ at both resolutions for both seen and unseen face swapping manipulations except for REFace. This implies that the matching rates between $m_{s}$ and $m_{\textrm{rec}}$ for images after common manipulations (real) are consistently higher than those for images after face swapping manipulations (fake). In other words, our identity perceptual watermarking framework not only robustly survives image manipulations, but also favorably distinguishes real and fake samples with respect to the changes in facial identities. Meanwhile, REFace has brought the most challenges to our watermarks, although the watermark robustness is proved to be good in Table II and Table III ($94.61\%$ and $94.83\%$). This reflects the potential doubt that REFace generates faces that still contain similar identities as the original ones, making higher similarity between $m_{s}$ and $m_{\textrm{rec}}$ [^5] Since this refers to unsatisfactory face swapping performance, there is no route to improve the proposed watermarks accordingly.. As a result, for a mixture of all eight Deepfake models, including seven unseen ones, our approach retains satisfactory detection performance at $95.39\%$ and $97.57\%$ at $128\times 128$ and $256\times 256$ resolutions, respectively, against both GAN-based and diffusion-based Deepfake face swapping models, demonstrating superior generalization ability.
TABLE VIII: Top-N source tracing accuracy of watermarks on CelebA-HQ.
Model Acc@5 Acc@3 Acc@1
SepMark [21] Ours SepMark [21] Ours SepMark [21] Ours
SimSwap [60] 0.25% 99.96% 0.10% 99.96% 0.03% 99.89%
InfoSwap [61] 5.68% 99.99% 4.08% 99.99% 2.27% 99.92%
UniFace [62] 16.02% 99.43% 13.71% 99.36% 9.38% 98.72%
E4S [63] 69.63% 99.43% 68.57% 99.04% 66.12% 98.08%
DiffSwap [64] 1.33% 99.75% 0.97% 99.67% 0.37% 99.31%
FuseAnyPart [65] 1.28% 99.99% 1.08% 99.96% 0.30% 99.87%
Face-Adapter [66] 1.85% 86.47% 1.48% 84.03% 0.37% 77.36%
REFace [67] 4.69% 99.79% 2.34% 99.72% 0.80% 99.33%
Average 12.59% 98.10% 11.54% 97.72% 9.96% 96.56%
@@
TABLE IX: Robustness ablation test on common manipulations for easier and harder parameter values that are unseen during training. The bit-wise watermark recovery accuracies are reported. Long common manipulation names are abbreviated.
@@
IV-F Source Tracing
In this paper, the proposed identity perceptual watermark aims to perform Deepfake detection and source tracing concurrently based on its robustness. Therefore, besides demonstrating promising watermark recovery accuracies and Deepfake detection AUC scores in early sections, in this section, we evaluated the watermark ability on source tracing. In particular, we treated each identity as a unique class and evaluated the top-5 (ACC@5), top-3 (ACC@3), and top-1 (ACC@1) accuracies on the recovered watermarks with respect to the images after face swapping. In other words, for each recovered identity perceptual watermark, we matched it back to the original facial identities and recorded the percentage of cases such that the correct identity lies in the top-5, top-3, and top-1 watermark similarities. For a fair comparison, we selected the state-of-the-art watermarking framework with the most competitive robustness in terms of watermark recovery accuracy, SepMark, as the contrastive method in the source tracing experiment.
The source tracing performance of the watermarks is reported in Table VIII. It can be observed that our approach promisingly preserves the top-5 accuracies above $99\%$ against all face swapping algorithms except for Face-Adapter. Meanwhile, even as the rule becomes more strict for top-3 and top-1 accuracies, the statistics are still maintained above $98\%$. As for the challenging Face-Adapter, the results are aligned with Table III where it causes a relatively lower watermark recovery accuracy of $87.23\%$. On the other hand, the promising source tracing performance against REFace further proves the fact that the Deepfake detection performance in Table VII is caused by the unsatisfactory generative performance of REFace. Contrarily, except performing $69.93\%$, $68.57\%$, and $66.12\%$ accuracies for top-5, top-3, and top-1 against E4S, the opponent that SepMark gains the best watermark robustness, SepMark derives generally poor accuracies below $20\%$ for all other face swapping models. As a result, our approach has been proved to have outstanding source tracing ability to find the original victim facial identities by providing reliable top-5, top-3, and top-1 accuracies regarding the face swapping manipulations.
TABLE X: Demonstration of the cyclic oscillations of $x_{i}$ with $x_{0}=0.1$ for different $r$ values.
$x_{0}=0.1$ $r=2$ $r=2.5$ $r=3.5$ $r=3.93$
$x_{1}$ 0.18 0.225 0.315 0.3537
$x_{2}$ 0.2952 0.435938 0.755213 0.898383
$x_{3}$ 0.416114 0.614740 0.647033 0.358772
$x_{4}$ 0.485926 0.592086 0.799334 0.904114
$x_{5}$ 0.499603 0.603800 0.561395 0.340696
$x_{10}$ 0.5 0.599880 0.864143 0.611306
$x_{11}$ 0.5 0.600060 0.410898 0.933811
$x_{12}$ 0.5 0.599970 0.847213 0.242905
$x_{13}$ 0.5 0.600015 0.453051 0.722736
$x_{14}$ 0.5 0.599992 0.867285 0.787528
$x_{15}$ 0.5 0.600003 0.402856 0.657598
$x_{16}$ 0.5 0.599998 0.841970 0.884890
$x_{17}$ 0.5 0.600001 0.465697 0.400310
$x_{18}$ 0.5 0.599999 0.870882 0.943443
$x_{19}$ 0.5 0.600000 0.393564 0.209698
$x_{20}$ 0.5 0.599999 0.835350 0.651298
IV-G Ablation Studies